OpenStudy (mayankdevnani):

A man of weight 60 kg.wt takes a body of mass 15 kg at a height of 10 m on a building in 3 minutes. The efficiency of man is - a) 10% b) 20% c)30% d) 40%

OpenStudy (mayankdevnani):

@experimentX @estudier @UnkleRhaukus @hartnn @.Sam. @phi @jhonyy9 @jiteshmeghwal9 @jiteshmeghwal9

OpenStudy (mayankdevnani):

@Jonask @Callisto @Hero @Yahoo! @hba

OpenStudy (hba):

The first step would be to make data :)

OpenStudy (mayankdevnani):

how?

OpenStudy (hba):

Data: W=60 kg h=10 m t=3 min--convert into seconds m=15 kg

OpenStudy (hba):

Do you know how to convert it into seconds ? @mayankdevnani

OpenStudy (anonymous):

times to 60 sekon., ok

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

t=3*60=180 s

OpenStudy (mayankdevnani):

right @hba

OpenStudy (hba):

Well and yeah efficiency is unknown so include it in the data :) \[\in=?\]

OpenStudy (hba):

Now See the appropriate formula,

OpenStudy (anonymous):

efisiensi = (using power/initial power) x 100%

OpenStudy (hba):

efficiency formula = (Wout/Win)*100%

OpenStudy (hba):

Wout is the work done by the machine is called the output work Win is the work done by you on a machine is called the input work efficiency= useful work output/work input

OpenStudy (mayankdevnani):

 i think it is a proper sign

OpenStudy (hba):

Yeah right :D

OpenStudy (anonymous):

yes

OpenStudy (hba):

Work=Force*distance

OpenStudy (hba):

P = w/t

OpenStudy (anonymous):

\[ \eta\]

OpenStudy (mayankdevnani):

hey!! symbol=? any name

OpenStudy (hba):

Efficiency

OpenStudy (mayankdevnani):

no it;'s eta

OpenStudy (hba):

Oh Eta

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

yes eta is eficiency symbol

OpenStudy (mayankdevnani):

then

OpenStudy (anonymous):

efficiency formula = (Pout/Pin)*100%

OpenStudy (mayankdevnani):

@TheViper

OpenStudy (mayankdevnani):

@experimentX

OpenStudy (theviper):

LOL you called every member !!!!!!!!

OpenStudy (mayankdevnani):

but i need solution

OpenStudy (mayankdevnani):

@Yahoo!

OpenStudy (mayankdevnani):

can anyone solve my problem......

OpenStudy (theviper):

I think everyone is trying :)

OpenStudy (mayankdevnani):

yaa... but no one is answering...

OpenStudy (mayankdevnani):

@Preetha

OpenStudy (anonymous):

@mayankdevnani is the answer 20%???

OpenStudy (mayankdevnani):

yaa but how? @Yahoo!

OpenStudy (anonymous):

efficiency formula = (Wout/Win)*100%

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

Wat is the Work Done by the Man... W = 15 *10 * 10=1500 (W=f*d)

OpenStudy (anonymous):

Work Done on the man is 60 +15 = 75 kg W = 75 *10*10 = 7500

OpenStudy (anonymous):

Eff = (1500/7500)*100 = 20%

OpenStudy (anonymous):

@mayankdevnani u got it.....)

OpenStudy (mayankdevnani):

1 min....

OpenStudy (mayankdevnani):

thnx....... @Yahoo!